OpenStudy (anonymous):

Use the function below to find F(-2). F(x) = 3x A. -9 B. -6 C. D.

OpenStudy (studygurl14):

Just plug in -2 into the function for x

OpenStudy (anonymous):

What is 3*-2?

OpenStudy (studygurl14):

So F(-2) = 3(-2) = ?

OpenStudy (anonymous):

-6

OpenStudy (studygurl14):

yep

OpenStudy (anonymous):

Exactly.

OpenStudy (michele_laino):

hint: we can write: \[\huge F\left( { - 2} \right) = 3 \cdot \left( { - 2} \right) = ...?\]